# Multiscale Dynamical Indices Reveal Scale-Dependent Atmospheric Dynamics

### v1

@@ 4.1 Scale-dependent circulation dynamics during heatwaves
In Fig. 2, we present the composite maps of the standardized anomalies of dynamical indices for the onset days of European summertime heatwaves (we will use heatwaves hereafter, for the sake of brevity, when referring to European summer heatwaves). We computed the dynamical indices using SLP (Fig. 2, rows 1 and 2) and Z500 (Fig. 2, rows 3 and 4) separately to show the distinct dynamical behavior of the surface-level and mid-level troposphere, following [15]. The composite maps of the standardized anomalies of the dynamical indices derived from the two observables reveal clear differences in both magnitude and spatial configuration, with the Z500-based indices generally showing stronger anomalies.
Figure 2: Composite of standardized anomalies of dynamical indices for the onset days of European summertime heatwaves. Composite of standardized anomalies for the local dimension $d$ (rows 1 and 3) and the inverse persistence $\theta$ (rows 2 and 4) that were computed using both SLP (rows 1 and 2) and Z500 (rows 3 and 4). Each column represents heatwaves over a specific region, with the affected area indicated by a colored solid bounding box. The dashed bounding boxes in the last row highlight regions utilized in the analysis presented in Fig. 4. The regions studied are as follows: British Isles (blue; solid box: 49°N–59°N, 10°W–2°E; dashed box: 36°N–62°N, 40°W–10°E), Central Europe (black; solid box: 45°N–55°N, 4°E–16°E; dashed box: 30°N–60°N, 30°W–20°E), Iberian Peninsula (orange; solid box: 36°N–44°N, 10°W–3°E; dashed box: 30°N–53°N, 38°W–5°E), Mediterranean (purple; solid box: 36°N–44°N, 10°E–25°E; dashed box: 28°N–54°N, 25°W–25°E), and Scandinavia (green; solid box: 57°N–65°N, 25°E–20°E; dashed box: 40°N–67°N, 30°W–24°E). Cross-hatching indicate statistical significance.
In the mid-level troposphere, significant negative $\theta$ anomalies (Fig. 2, row 4) are observed for Z500 slightly west of the target regions. This suggests that during the onset days of heatwaves, the upstream atmospheric configuration of the affected regions is unusually persistent, but other regions in the Euro-Atlantic sector do not have increased persistence. Indeed, less persistent atmospheric configurations with relatively smaller spatial extent can be found over the Atlantic Ocean for all studied regions (except Iberia). We note that the results presented in Fig. 2 are robust with respect to the sliding window size, as similar spatial anomaly patterns are observed when a smaller sliding window is applied (see SI Fig. S1).
@@
In addition, the local dimension ($d$) of Z500 (Fig. 1, row 3) shows some negative anomalies that are co-located with, but weaker than, the negative anomalies of inverse persistence ($\theta$). Other regions display relatively weak or near-zero anomalies, except for British Isles heatwaves, that exhibit a strong positive anomaly over the Atlantic. Such results suggest that Z500 patterns display reduced dynamical complexity during heatwave onset days near the affected regions. The results for $\theta$ and $d$ on Z500 together support the view that European summertime heatwaves are accompanied by persistent and low-dimensional atmospheric configurations in the mid-troposphere from a dynamical systems perspective.
@@
At the surface level, we note that the standardized anomalies are less pronounced, yet still reveal some coherent patterns (Fig. 2). For heatwaves occurring outside Iberia, we observe significant negative anomalies in both local dimension ($d$) and inverse persistence ($\theta$) near the affected regions, similar to those in the mid-level atmosphere, although much weaker. In Iberia, no such significant anomalies are observed for $\theta$ over the entire domain. It is particularly noteworthy to observe distinct dynamical properties at the surface and mid-level troposphere, as also noted by [15], especially since these differences are not evident in the composite anomaly maps of the raw variables (see SI Fig. S2). This is also consistent with meteorological perspectives, as heat lows can develop during heatwaves [53], exhibiting dynamics that differ significantly from the mid-tropospheric flow. At the same time, previous studies have shown that blocking events, which are frequently associated with heatwaves, are primarily reflected at the mid- or upper-level flows [54, 55].
@@
Considering the importance of blocking and the non-systematic link between heatwaves and blocking, we further extend our analysis to blocking events to enhance the completeness of this study, particularly given previously noted contrasting perspectives on their dynamical properties. From a meteorological perspective, blocking events are persistent atmospheric configurations by definition [43, 54, 55]. However, recent studies grounded in dynamical systems theory propose a differing view, suggesting that blocking regimes are not inherently persistent atmospheric configurations [13, 17]. In order to reconcile differing these views, we apply a blocking detection algorithm to examine the dynamical properties of blocking following [56], with detailed steps provided in SI Text S3. We find that, regardless of their location, blocking events are associated with geographically localized more persistent (low $\theta$) and less complex (low $d$) mid-level tropospheric flows – see SI Fig. S5, where the second and third rows represent the anomalies of $d$ and $\theta$ for the blocking events identified by the blocking algorithm adopted. Yet, regions surrounding the blocking center are less persistent (high $\theta$) and more complex (high $d$), and may be less predictable as also shown in operational forecasts [55], and in agreement with the theoretical arguments set in [47]. This further analysis on blocking, bridges the two seemingly different views that were present in the literature, thereby showing how the new approach can capture relevant informative features of atmospheric dynamics.
@@
A further fascinating aspect emerges when looking at the temporal coherence of the anomaly fields of the dynamical indicators (as well as the meteorological fields themselves) as we approach the onset of the heatwaves and as we observe its decay, see Figs. S3-S5 in the SI. In [57, 58] it was argued, on the basis of large deviation theory, that persistent extreme events like heatwaves and cold spells are characterized by dynamical typicality. This means that the occurrence of the extreme event or interest - e.g. heatwave in the British isles - requires a very special large-scale configuration of the atmospheric fields, which to a very good degree of approximation repeats itself, thus defining a class of analogues, for different individual occurrences of the event. In other terms, Such a configuration is atypical with respect to the overall statistics of the fields, but becomes typical when we focus on the days associated with the heatwave of interest. Additionally, it was argued that special large-scale structures coherently evolve in the period leading to the onset of the extreme event. A separate confirmation of this idea was given on the 2021 Pacific Northwest heatwave by [59]. Indeed, we find that in the days preceding the onset of the heatwaves the composite fields have a rather high degree of coherence, providing good evidence for the existence of a highway in phase space (technically associated with an instantonic trajectory) leading to the events; see discussion in [60] and first original contribution is this direction in [61]. Note that such coherence is largely lost after the decay of the heatwaves, because the lysis process, as opposed to the onset, does not follow a preferential path.
Figure 3: Box plots of standardized anomalies of dynamical indices for the Euro-Atlantic sector using the single-domain approach. The dynamical indices are computed within the large-scale bounding box (20°N–80°N, 80°W–40°E), followed by standardization steps as described above. Each subplot shows the results for one studied region, with statistically significant results highlighted in the color representing that region, as in Fig. 1. The ends of the boxes indicate the 25th and 75th percentiles, while the whiskers extend to 1.5 times the interquartile range beyond the box.
@@ 4.2 A richer dynamical view
Figure 4: Box plots of standardized anomalies of dynamical indices for the dashed bounding boxes shown in Fig. 2 using the single-domain approach. The dynamical indices are computed within the the dashed bounding boxes shown in the row 4 of Fig. 2, followed by standardization steps as described above. Each subplot shows the results for one studied region, with statistically significant results highlighted in the color representing that region, as in Fig. 1. The ends of the boxes indicate the 25th and 75th percentiles, while the whiskers extend to 1.5 times the interquartile range beyond the box.
@@
To further complement our analysis with the new framework, we conducted a comparative study following the previous single-domain approach – treating the entire Euro-Atlantic domain as a single dynamical system. Fig. 3 presents box plots of standardized anomalies of dynamical indices for the entire Euro-Atlantic domain during the onset days of summertime heatwaves, with each subplot corresponding to a specific column (i.e., affected region considered) in Fig. 2. Unlike our new approach in Fig. 2, that shows significant anomalies for both $d$ and $\theta$ in all regions considered, the single-domain approach shows that significant results are achieved in only three cases (out of twenty). This suggests that atmospheric patterns associated with heatwaves exhibit a range of different dynamical properties, that are difficult to capture via the single-domain approach. Specifically, using the single-domain method, at the mid-level, both the British Isles and Central Europe exhibit near-zero standardized anomalies, whereas the other three regions display negative anomalies for both indices. However, only the inverse persistence ($\theta$) for the Iberian Peninsula and the Mediterranean is statistically significant. At the surface level, no significant results are observed, except for the SLP-based inverse persistence ($\theta$), which shows positive standardized anomaly over the British Isles.
@@
These results are consistent with [15], showing that summertime heatwaves in Europe are not systematically linked with more persistent large-scale circulation patterns. We do not consider this to be in conflict with the findings from our new approach (Fig. 2); rather, it underscores the importance of incorporating different spatial scales in the analysis. Our new approach can pinpoint the regions that exhibit the most dynamical relevance of the studied phenomena, a result that was not possible to achieve with the pre-selected fixed-domain approach. Indeed, based on the results shown in Fig. 2, we can identify bounding boxes where the mid-level atmosphere demonstrates significantly higher persistence (dashed bounding boxes in row 4, Fig. 2). Different bounding boxes are used for heatwaves in different affected regions, as expected due to the assumption that events in different regions may have diverse atmospheric drivers. We compute the dynamical indices of atmospheric circulation patterns within each bounding box, and plot the standardized anomalies as box plots for the heatwave onset days in the corresponding regions (see Fig. 4). Consistent with expectations, more statistically significant results are observed when smaller bounding boxes are used. In particular, significantly more persistent characteristics are displayed at the mid-level troposphere during the onset days of heatwaves for all regions. Similarly, negative anomalies in local dimension ($d$) are observed across all regions, although they are statistically significant only in the Mediterranean and Scandinavia. At the surface level, only the local dimension ($d$) in Scandinavia exhibits a significant negative anomaly, while the others are not significant.
@@ 5 Discussions and conclusions
Dynamical systems theory has been extensively applied to diverse complex physical systems across various spatial scales [33, and references therein]. However, when addressing spatio-temporal systems, past studies have computed the metrics on a predefined bounding box based on prior expert knowledge. While sensitivity analyses on the boundaries of such bounding boxes are often discussed, the influence of the bounding box’s size – namely, the considered spatial scale – is rarely examined. To address these constraints, we propose a novel framework for applying dynamical indices to spatio-temporal complex systems, inspired by the concept of convolutional neural networks that leverage localized filters to extract spatially and temporally relevant features. By employing a sliding bounding box with a customizable size, we extract the system’s dynamical properties that are ‘local’ in both time (i.e., instantaneous) and space (i.e., spatially specific) at the spatial scale of interest. Compared to the previous approach with only two indices as outputs, this method generates spatio-temporal maps that allow for a more detailed analysis, providing richer insights into the dynamical features of certain phenomena of interest.
@@
In this work, we applied this new framework to high-impact European summertime heatwaves to demonstrate its effectiveness. Our results offer an explanation that reconciles the previously noted discrepancy in the persistence of heatwaves-associated circulation patterns between the meteorological view and the dynamical systems theory perspective [15]. In fact, the conclusions drawn from both perspectives are valid within their respective contexts, where the glue between the two is provided by the varying dynamical characteristics of heatwaves across different spatial scales. Although heatwaves are not systematically linked to large-scale circulation patterns with distinct dynamical properties, they often display significant persistence and low-dimensional signatures in the regions upstream of the affected areas. Additionally, looking at the temporal evolution of the anomaly fields of the dynamical indicators and of the meteorological fields themselves we find further confirmation of the validity of the notion of dynamical typicality of persistent extreme events discussed in [57, 58].
@@
More broadly, we extended our analysis to summertime blocking events that frequently cause heatwaves in Europe, a phenomenon considered highly persistent by definition but, according to some recent studies employing various methodologies – including dynamical systems theory [13, 46], mathematical models [47], and ensemble forecasts [62] – is regarded as a relatively unpredictable configuration. Similar to the findings for heatwaves, we observed that blocking centers exhibit significantly greater persistence and predictability, while being accompanied by configurations with contrasting dynamical properties in other regions within the Euro-Atlantic sector. This indicates that the dynamical properties of blocking are closely tied to spatial scales, and analyses at different spatial scales have resulted in divergent conclusions in previous studies [45, 13]. Thus, we believe that our approach bridges the perspective of dynamical systems theory and more conventional meteorological views, providing an explanation for the divergent result present in the literature.
@@
As a novel framework, we anticipate its broad application in future research rooted in dynamical systems theory, as it is general in nature and can be applied to any region or variable in atmospheric dynamics, as well as other geophysical systems (e.g., slow earthquakes [25]). For instance, it could be utilized to study energy cascades [63] in turbulent flow to better understand the transfer of energy across different scales. In atmospheric science, we believe it has the potential to open up research opportunities for a deeper understanding of teleconnection dynamics [64], as our analysis provides spatio-temporal dynamical information. At the same time, the temporal evolution of dynamical indices, which has been studied in the life cycles of weather regimes [18, 17] and extreme events [14], can be further complemented by our framework to provide additional insights. Moreover, it could potentially be used to identify distant dynamical precursors of the studied phenomena – another aspect that cannot be achieved using the fixed-domain approach. This holds great significance for advancing our understanding of their onset mechanisms and potentially enhancing predictive modeling.
Acknowledgments
G.M. acknowledges support from MOE Tier 2 grant no 22-5191-A0001-0, titled: ‘Prediction-to-Mitigation with Digital Twins of the Earth’s Weather’. V.L. acknowledges the support provided by the Horizon Europe Project ClimTIP (Grant No. 100018693) and by the EPSRC project LINK (Grant No. EP/Y026675/1).
Supplementary Information
Text S1. Dynamical Indices

### v2

@@ 4.1 Scale-dependent circulation dynamics during heatwaves
In Fig. 2, we present the composite maps of the standardized anomalies of dynamical indices for the onset days of European summertime heatwaves (we will use heatwaves hereafter, for the sake of brevity, when referring to European summer heatwaves). We computed the dynamical indices using SLP (Fig. 2, rows 1 and 2) and Z500 (Fig. 2, rows 3 and 4) separately to show the distinct dynamical behavior of the surface-level and mid-level troposphere. The composite maps of the standardized anomalies of the dynamical indices derived from the two observables reveal clear differences in both magnitude and spatial configuration, with the Z500-based indices generally showing stronger anomalies.
@@
In the mid-level troposphere, significant negative $\theta$ anomalies (Fig. 2, row 4) are observed for Z500 slightly west of the target regions. This suggests that during the onset days of heatwaves, the upstream atmospheric configuration of the affected regions is unusually persistent, but other regions in the Euro-Atlantic sector do not have increased persistence. In addition, the local dimension ($d$) of Z500 (Fig. 1, row 3) shows some negative anomalies that are co-located with, but weaker than, the negative anomalies of inverse persistence ($\theta$). Other regions display relatively weak or near-zero anomalies, except for British Isles heatwaves, that exhibit a strong positive anomaly over the Atlantic. Such results suggest that Z500 patterns display reduced dynamical complexity during heatwave onset days near the affected regions. The results for $\theta$ and $d$ on Z500 together support the view that European summertime heatwaves are accompanied by persistent and low-dimensional atmospheric configurations in the mid-troposphere from a dynamical systems perspective. We note that the results presented in Fig. 2 are robust with respect to the sliding window size, as similar spatial anomaly patterns are observed when a smaller sliding window is applied (see Fig. 6).
Figure 2: Composite of standardized anomalies of dynamical indices for the onset days of European summertime heatwaves. Composite of standardized anomalies for the local dimension $d$ (rows 1 and 3) and the inverse persistence $\theta$ (rows 2 and 4) that were computed using both SLP (rows 1 and 2) and Z500 (rows 3 and 4). Each column represents heatwaves over a specific region, with the affected area indicated by a colored solid bounding box. The dashed bounding boxes in the last row highlight regions utilized in the analysis presented in Fig. 5. The regions studied are as follows: British Isles (blue; solid: 49°N–59°N, 10°W–2°E; dashed: 36°N–62°N, 40°W–10°E), Central Europe (black; solid: 45°N–55°N, 4°E–16°E; dashed: 30°N–60°N, 30°W–20°E), Iberian Peninsula (orange; solid: 36°N–44°N, 10°W–3°E; dashed: 30°N–53°N, 38°W–5°E), Mediterranean (purple; solid: 36°N–44°N, 10°E–25°E; dashed: 28°N–54°N, 25°W–25°E), and Scandinavia (green; solid: 57°N–65°N, 25°E–20°E; dashed: 40°N–67°N, 30°W–24°E). Cross-hatching indicate statistical significance.
@@
At the surface level, we note that the standardized anomalies are less pronounced, yet still reveal some coherent patterns (Fig. 2). For heatwaves occurring outside Iberia, we observe significant negative anomalies in both local dimension ($d$) and inverse persistence ($\theta$) near the affected regions, similar to those in the mid-level atmosphere, although much weaker. In Iberia, no such significant anomalies are observed for $\theta$ over the entire domain. It is particularly noteworthy to observe distinct dynamical properties at the surface and mid-level troposphere, as also noted by Holmberg et al. [10], especially since these differences are not evident in the composite anomaly maps of the raw variables (see Fig. 7). Such contrasts are consistent with meteorological perspectives, as surface and mid-tropospheric dynamics can differ substantially, with many features relevant to surface weather manifesting at mid-levels. This further illustrates that the SDI framework, through its broad applicability and sensitivity to variable choice, can reveal distinct and complementary dynamical information, thereby serving as a useful diagnostic tool for complex systems.
Considering the importance of blocking and the non-systematic link between heatwaves and blocking, we further extend our analysis to blocking events to enhance the completeness of this study, particularly given previously noted contrasting perspectives on their dynamical properties. From a meteorological perspective, blocking events are persistent atmospheric configurations by definition, blocking events are by definition persistent atmospheric configurations as their identification involves applying a minimum duration threshold to a particular index [35, 51, 52]. However, recent studies grounded in dynamical systems theory propose a differing view, suggesting that blocking regimes are not inherently persistent atmospheric configurations [8, 12].
@@
In order to reconcile these views, we apply a blocking detection algorithm to examine the dynamical properties of blocking [53], with detailed steps provided in Appendix C. We find that, regardless of their location, blocking events are associated with geographically localized more persistent (low $\theta$) and less complex (low $d$) mid-level tropospheric flows – see Fig. 9, where the second and third rows represent the anomalies of $d$ and $\theta$ for the blocking events identified by the blocking algorithm adopted. Yet, regions surrounding the blocking center are less persistent (high $\theta$) and more complex (high $d$), and may be less predictable as also shown in operational forecasts [52] and in agreement with the theoretical arguments set in Schubert et al [40]. This means while predictability increases within the blocking region, this improvement is more than offset by reduced predictability in the surrounding areas, leading to reduced predictability from a large-scale perspective. This further analysis of blocking bridges the two seemingly different views that were present in the literature, thereby showing how the new approach can capture relevant informative features of atmospheric dynamics.
Figure 3: Temporal evolution of the composite anomaly of Z500-based inverse persistence ($\theta$) before the onset and after the decay of European summertime heatwaves. Composite anomaly of Z500-based inverse persistence from two days before the onset to the onset day, and from the last day to two days after.
@@
A further fascinating aspect emerges when looking at the temporal coherence of the anomaly fields of the dynamical indicators as we approach the onset of the heatwaves and as we observe its decay. In this context, Figure 3 presents the composite anomaly maps of the Z500-based inverse persistence field. This variable was chosen for its strong dynamical relevance to heatwaves, as demonstrated by the results in Fig. 2, while the evolution of the Z500-based local dimension is shown in Fig. 8. We observed that the onset of heatwaves exhibits a higher degree of coherence than the decay process, despite some variability across affected regions. Previous studies, based on large deviation theory, have argued that persistent extreme events such as heatwaves and cold spells exhibit dynamical typicality [44, 45]. This means that the occurrence of the extreme event requires a very special large-scale configuration of the atmospheric fields, which to a very good degree of approximation repeats itself, thus defining a class of analogues, for different individual occurrences of the event. In other terms, such a configuration is atypical with respect to the overall statistics of the fields, but becomes typical when we focus on the days associated with the heatwave of interest. Indeed, our findings provide good evidence for the existence of a highway in phase space leading to the events; see discussion in Galfi et al. [54] and first original contribution is this direction by Dematteis et al. [55]. Such coherence is largely lost after the decay of the heatwaves, because the lysis process, as opposed to the onset, does not follow a preferential path. Our analysis could also be used to identify small-scale dynamical precursors of heatwaves – or other extreme events in general. However, this would require a more detailed clustering or case study of these events, accounting for the diverse dynamical pathways leading to them [23], which we consider beyond the scope of this methodological paper.
@@ 4.2 A richer dynamical view
Figure 4: Box plots of standardized anomalies of dynamical indices for the Euro-Atlantic sector using the single-domain approach. The dynamical indices are computed within the large-scale bounding box (20°N–80°N, 80°W–40°E), followed by standardization steps as described above. Each subplot shows the results for one studied region, with statistically significant results highlighted in the color representing that region, as in Fig. 1.
@@
To further complement our analysis with the new framework, we conducted a comparative study following the previous single-domain approach – treating the entire Euro-Atlantic sector as a single dynamical system. Fig. 4 presents box plots of standardized anomalies of dynamical indices for the entire Euro-Atlantic domain during the onset days of summertime heatwaves, with each subplot corresponding to a specific column (i.e., affected region considered) in Fig. 2. Unlike our new approach in Fig. 2, that shows significant anomalies for both $d$ and $\theta$ in all regions considered, the single-domain approach shows that significant results are achieved in only three cases. This suggests that atmospheric patterns associated with heatwaves fit a range of different dynamical properties, which are difficult to capture via the single-domain approach. Specifically, using the single-domain method, at the mid-level, both the British Isles and Central Europe exhibit near-zero standardized anomalies, whereas the other three regions display negative anomalies for both indices. However, only the inverse persistence ($\theta$) for the Iberian Peninsula and the Mediterranean is statistically significant. At the surface level, no significant results are observed, except for the SLP-based inverse persistence ($\theta$), which shows positive standardized anomaly over the British Isles.
Figure 5: Box plots of standardized anomalies of dynamical indices for the dashed bounding boxes shown in Fig. 2 using the single-domain approach. The dynamical indices are computed within the dashed bounding boxes shown in the row 4 of Fig. 2, followed by standardization steps as described above. Each subplot shows the results for one studied region, with statistically significant results highlighted in the color representing that region, as in Fig. 1.
@@
These results are consistent with Holmberg et al.[10], showing that summertime heatwaves in Europe are not systematically linked with more persistent large-scale circulation patterns. We do not consider this to be in conflict with the findings from our new approach (Fig. 2); rather, it underscores the importance of incorporating different spatial scales in the analysis. Our new approach can pinpoint the regions that exhibit the most dynamical relevance of the studied phenomena, a result that was not possible to achieve with the pre-selected fixed-domain approach. Indeed, based on the results shown in Fig. 2, we can identify bounding boxes where the mid-level atmosphere demonstrates significantly higher persistence (dashed bounding boxes in row 4, Fig. 2). Different bounding boxes are used for heatwaves in different affected regions, as expected due to the assumption that events in different regions may have diverse atmospheric drivers. We compute the dynamical indices of atmospheric circulation patterns within each bounding box, and plot the standardized anomalies as box plots for the heatwave onset days in the corresponding regions (see Fig. 5). Consistent with expectations, more statistically significant results are observed when smaller bounding boxes are used. In particular, significantly more persistent characteristics are displayed at the mid-level troposphere during the onset days of heatwaves for all regions. Similarly, negative anomalies in local dimension ($d$) are observed across all regions, although they are statistically significant only in the Mediterranean and Scandinavia. At the surface level, only the local dimension ($d$) in Scandinavia exhibits a significant negative anomaly, while the others are not significant.
@@ 5 Discussions and conclusions
Dynamical systems theory has been extensively applied to diverse complex physical systems across various spatial scales [28]. However, when addressing spatio-temporal systems, past studies have computed the metrics on a predefined bounding box based on expert knowledge. While sensitivity analyses on the boundaries of such bounding boxes are often discussed, the influence of the bounding box’s size – namely, the considered spatial scale – is rarely examined. To address these constraints, we propose a novel framework for applying dynamical indices to spatio-temporal complex systems, inspired by the concept of convolutional neural networks. By employing a sliding bounding box with a customizable size, we extract the system’s dynamical properties that are ‘local’ in both time and space at the spatial scale of interest. Compared to the previous approach with only two indices as outputs, this method generates spatio-temporal maps that allow for a more detailed analysis, providing richer insights into the dynamical features of certain phenomena of interest.
Despite these advantages, certain caveats of this framework should be acknowledged. Notably, it is significantly more computationally expensive than previous approach, and this cost cannot be mitigated. Additionally, the choice of sliding window size relies on prior knowledge of the system. In many cases, obtaining meaningful insights requires analyzing results from different sliding window sizes, further increasing computational costs. To address this, it is advisable to start with a relatively coarse version of SDIs (i.e., a large stride size) to manage computational costs, while noting that the computation can be parallelized to further enhance efficiency. Meanwhile, our framework inherits an inherent limitation of dynamical indices: their computation relies on assumptions that the available data may not fully satisfy, and their impact cannot be quantified [56].
@@
In this work, we applied this new framework to high-impact European summertime heatwaves to demonstrate its effectiveness. Our results offer an explanation that reconciles the previously noted discrepancy in the persistence of heatwaves-associated circulation patterns between the meteorological view and the dynamical systems theory perspective [10]. In fact, the conclusions drawn from both perspectives are valid within their respective contexts, where the glue between the two is provided by the varying dynamical characteristics of heatwaves across different spatial scales. Although heatwaves are not systematically linked to large-scale circulation patterns with distinct dynamical properties, they often display significant persistence and low-dimensional signatures in the regions upstream of the affected areas. Additionally, looking at the temporal evolution of the anomaly fields of the dynamical indicators and of the meteorological fields themselves we find further confirmation of the validity of the notion of dynamical typicality of extreme events [44, 45].
@@
More broadly, we extended our analysis to summertime blocking events that frequently cause heatwaves in Europe, a phenomenon considered highly persistent by definition but, according to some recent studies employing various methodologies – including dynamical systems theory [8, 38], mathematical models [39], and ensemble forecasts [57] – is regarded as a relatively unpredictable configuration. Similar to the findings for heatwaves, we observed that blocking centers exhibit significantly greater persistence and predictability, while being accompanied by configurations with contrasting dynamical properties in other regions within the Euro-Atlantic sector (see also Springer et al. [17]). This indicates that the dynamical properties of blocking are closely tied to spatial scales, and analyses at different spatial scales have resulted in divergent conclusions in previous studies [37, 8]. Thus, we believe that our approach bridges the perspective of dynamical systems theory and more conventional meteorological views, providing an explanation for the divergent result present in the literature.
@@
As a novel framework, we anticipate its broad application beyond the studied variables and domain, and even beyond atmospheric systems, to other spatio-temporal complex systems where spatial scales play a critical role. For instance, it could be utilized to study energy cascades [58] in turbulent flow to better understand the transfer of energy across different scales. In atmospheric science, its general nature allows it to be applied to any region or variable in atmospheric dynamics, particularly in the tropics, where the interplay of complex internal variabilities makes scale considerations especially crucial for analysis [59]. Meanwhile, the temporal evolution of dynamical indices, which has been studied in the life cycles of weather regimes [13, 12] and extreme events [9], can be further complemented by our framework to provide additional insights. Moreover, it could potentially be used to identify distant dynamical precursors of the studied phenomena – another aspect that cannot be achieved using the fixed-domain approach. This holds great significance for advancing our understanding of their onset mechanisms and potentially enhancing predictive modeling.
Figure 8: Temporal evolution of the composite anomaly of Z500-based local dimension ($d$) before the onset and after the decay of European summertime heatwaves. As in Fig. 3, but for Z500-based local dimension.
Appendix A Dynamical Indices